{F} 2. Find the particular solution of the differential equation 1 − cos(2t) dx/dt = x sin(2t) given that x(π/2) = 2.

1
Expert's answer
2022-02-16T09:04:30-0500

Let us find the particular solution of the differential equation "(1 − \cos(2t))\frac{ dx}{dt} = x \sin(2t)" given that "x(\frac{π}2) = 2."

This equation is equivalent to "\frac{ dx}{x} = \frac{\sin(2t)dt}{1 − \cos(2t)}", and hence "\int\frac{ dx}{x} = \int\frac{\sin(2t)dt}{1 − \cos(2t)}" .

It follows that

"\ln|x|= \frac{1}2\int\frac{d(1 − \cos(2t))}{1 − \cos(2t)}=\frac{1}2\ln|1 − \cos(2t)|+\ln|C|=\ln|C\sqrt{1 − \cos(2t)}|."


We conclude that "x(t)=C\sqrt{1 − \cos(2t)}" is the general solution.


If "x(\frac{π}2) = 2" then "2=C\sqrt{1-\cos\pi}=C\sqrt{2}." It follows that "C=\sqrt{2}," and thus the particular solution is

"x(t)=\sqrt{2}\sqrt{1 − \cos(2t)}=\sqrt{2 − 2\cos(2t)}."